EPS Junkyard Arts, Inc., had earnings of $161,700 for the year. The company had 28,000 shares...

  1. EPS

    Junkyard Arts, Inc., had earnings of $161,700 for the year. The company had 28,000 shares of common stock outstanding during the year and issued 2,100 shares of $50 par value preferred stock. The preferred stock has a dividend of $5 per share. There were no transactions in either common or preferred stock during the year.

    Determine the basic earnings per share for Junkyard Arts for the year. Round answer to two decimal places.
    $per share
Answer #1

Preferred dividend = 2100*5 = 10500

Basic earning per share = (Net income-Preferred dividend)/Share outstanding = (161700-10500)/28000 = 5.40 per share